ganeshie8 (ganeshie8):

Mary wants to grow rose plants in her backyard. So she goes to the local nursery and spends a total of $4000 for 100 rose plants of 3 different colors : pink : $120 each white : $50 each red : $25 each Whas are the possible counts of each color plant ? Answer : (5, 41, 54) (10, 22, 68) (15, 3, 82)

OpenStudy (anonymous):

I think we can set up a system of equations for this.

OpenStudy (ikram002p):

mmm 120 r +50 w+ 25 re =4000 r+w+re =100 and anyother condition ?

OpenStudy (anonymous):

\[120p+50w+25r = 4000\] \[p + w + r = 100\]

OpenStudy (anonymous):

3 Variables 2 equations!?

ganeshie8 (ganeshie8):

yes, that was all the given problem... and I have the answer if it helps..

OpenStudy (ikram002p):

its not a probelm , its OR problem u can add other condition which is all r,w,re >0 ans by graph get the integer solution

ganeshie8 (ganeshie8):

yes p,w,r have to be whole numbers as they represent the number of plants

OpenStudy (kainui):

I had half a rose before

ganeshie8 (ganeshie8):

*natural numbers

OpenStudy (ikram002p):

wait i have QSB lets see if it could solve it =)

OpenStudy (anonymous):

Should we set up another equation using the 3 colors = ?

ganeshie8 (ganeshie8):

lol Mary wont buy half/quarter plants, let me add this constraint : she knows only natural numbers :)

ganeshie8 (ganeshie8):

another equation ?

ganeshie8 (ganeshie8):

I've modified the question slightly..

OpenStudy (kainui):

dividing through by 5 and eliminating a variable we get: 300=19p+5w since 300 is divisible by 5, and apparently since there's only one answer, she had to have gotten 0 pink flowers. 300=5w w=60 That means r=40 Check it. =)

OpenStudy (ikram002p):

how to graph two function in wolfram ?

OpenStudy (kainui):

So apparently "At least 1 flower" is something I didn't consider for my answer to work. Hmm. Now that I see the updated answers, I see I'm wrong.

ganeshie8 (ganeshie8):

120p + 50w + 25r = 4000 p+w+r = 100 r = 100 - p - w

ganeshie8 (ganeshie8):

24p + 10w + 5(100-p-w) = 800 19p + 5w = 300

ganeshie8 (ganeshie8):

yeah if we allow 0, we will get w=60, p = 0, r = 40 as one solution

OpenStudy (ikram002p):

ok so if u consider the line (changed variable to feel comfortable =D ) 19x+5y=300 can we use number theory xD i mean we wanna x,y integers

OpenStudy (kainui):

I think in a sense allowing any amount of flowers to be zero, we break it. We could buy all red flowers if we wanted.

ganeshie8 (ganeshie8):

true ! but that may not be possible since 120 does not go evenly in 4000

OpenStudy (kainui):

We could easily buy almost any combination of red and white flowers.

OpenStudy (ikram002p):

ok so since we need integer 16x+5y=300 x should devide 5 right , since we need integer ( remember GCD =) ) so x= 5 k or something

OpenStudy (kainui):

Yeah I think ikram is on the right track here. I am bad at number theory but this looks like a GCD sort of diophantine euclid's algorithm whatever idk.

ganeshie8 (ganeshie8):

ikram yes :) using number theory is fine, but please explain the solution :)

OpenStudy (ikram002p):

ok do u got the part that x have to be 5k ?

ganeshie8 (ganeshie8):

typo in your equation : 19x + 5y = 300

ganeshie8 (ganeshie8):

19x = 300-5y 19x = 5(60-y) since 5 cannot divie 19, 5 has to divide x so x = 5k fine till here :)

OpenStudy (ikram002p):

ok so x=5k * 19k +y=60 y=60-19k * so what integer k that give integer y xD ( dumb guess )

ganeshie8 (ganeshie8):

k = 0, 1,2 will do ?

OpenStudy (ikram002p):

yeah from 0 up to 3 gives positive

OpenStudy (ikram002p):

so we have k=0 y=60 k=1 y=41 k=2 y=22 k=3 y=3

ganeshie8 (ganeshie8):

wow this looks real simple now xD why do we eliminate k=0 case ?

OpenStudy (ikram002p):

so we have set of solution for x,y (0,60) (5,41 ) (10,22 ) (15 ,3 ) mm does that work or we need to check something :-\

ganeshie8 (ganeshie8):

Gotcha ! k=0 makes x also 0, so we eliminate k=0 case !!

OpenStudy (ikram002p):

well yeah , but still its solution though

ganeshie8 (ganeshie8):

haha nope, 0 is not natural ! it was brilliant work from everyone, thanks a lot =))

OpenStudy (neer2890):

why solving like this. we just have to find the answer x+y+z=100 120x+50y+25z=4000 try all x ,y and z to satisfy 1st and 2nd equation. for e.g., A:(5,41,54) so 1:5+41+54=100 2:120*5+50*41+25*54=4000 so 1st is correct.

ganeshie8 (ganeshie8):

to me, the solution we have worked here looks better than any other fancy number theoretic solutions using diophantine equations xD

OpenStudy (ikram002p):

haha well , what we did was diphontine =) (some how )

OpenStudy (ikram002p):

speeks of NT , im trying to find why a prime number N of all 1's digits are rare :3

ganeshie8 (ganeshie8):

I think i came across that topic before, you're referring to prime numbers of form `111111111...` right ?

OpenStudy (ikram002p):

yep

OpenStudy (ikram002p):

wait ill show u something

ganeshie8 (ganeshie8):

@neer2890 that works if we have the options already :)

ganeshie8 (ganeshie8):

yes

OpenStudy (neer2890):

but you have given all the options at first time, so i thought options are already available.

OpenStudy (anonymous):

lol,u gave the options,mee to thought first one,then by solving all options gives 4000